OpenStudy (anonymous):

given that the city c (population 702,000) is 700 miles from city d (population 550,000) what is the average number of daily phone calls between city c and city d

OpenStudy (tkhunny):

Is there a known relationship between population (direct) and distance (indirect)? Please show your work.

OpenStudy (anonymous):

oh sorry. first question was Given that the distance between city a (population 777,000) and city b (population 3695000) is 420 and that the average number fo daily phone calls between the cities is 427,000 what is the value of k? And I got an answer of 0.03. I am to use that answer for the new question.

OpenStudy (tkhunny):

You didn't answer my question. What are the theoretical relationships? \(Phone\;Calls = k\cdot\dfrac{Total\;Population}{Distance}\) I'm just guessing.

OpenStudy (anonymous):

sorry for the delay. im new to this forum just learning it . The equation used was c=\[c=\frac{ k*P _{1}*P _{2} }{ d ^{2} }\] using c= 427,000 k*777,000 * 3695000/420\[^{2}\]

OpenStudy (anonymous):

\[C=\frac{ k*777,000*3,695,000 }{ 420^{2} }\]

OpenStudy (tkhunny):

Lisa täyty mitä? With values this large you managed 0.03? I'm tempted to use WAY more decimal places and round the answer if you really aren't interested in that much precision. k = 0.0262356 The, we can solve it. \(0.0262356\dfrac{702000\cdot 550000}{700^{2}} = 20673\) or about 21000. Notice how k = 0.03 over estimates quite a bit. \(0.0262356\dfrac{702000\cdot 550000}{700^{2}} = 23639\) or about 24000. That's a whopping 14% over-estimate. You will spend too many resources to implement this plan.

OpenStudy (anonymous):

oh I agree. and thank you. My head is spinning right now.